gcc8617:

how much is the ratio 6:8 multiplied

TogaHimiko:

6:8 multiplied would be"6/8 = 3 x 2/4 x 2 = 3/4 "

TogaHimiko:

so your answer is 3/4

byakuyatogami1234:

here is the EQUASION ONLY SO DONT CHEW MY HEAD OFF!!! 6/8 = 3 x 2/4 x 2 = find it

gcc8617:

@byakuyatogami1234 wrote:
here is the EQUASION ONLY SO DONT CHEW MY HEAD OFF!!! 6/8 = 3 x 2/4 x 2 = find it
thx